On earth, two parts of a space probe weigh 18300 N and 4370 N. These parts are separated by a center-to-center distance of 7.71 m and may be treated as uniform spherical objects. Find the magnitude of the gravitational force that each part exerts on the other out in space, far from any other objects.
Answer #1

As these bodies are far in out space so no other force will act on them, there is only force, that act between them.

We know the formula,
F=(Gm1*m2)/d^2
where F is force
G is gravitational constant,G=6.67 x 10^-11 N m^2/kg^2
m1and m2 are the respective masses of thebodies
d is the distance

To get m1 and m2, divide the weight by9.8 m/s^2:
m1= 18300 N / 9.81 m/s^2 = 1865.44 kg
m2= 4370 / 9.81 = 445.46 kg

Now plug in the values to find the gravitational force:
F= (6.67 x 10^ -11 N-m^2/kg^2)(1865.44kg)(445.46 kg) / (7.71m)^2

=9.324x 10^-7N